# AutoMIR: Effective Zero-Shot Medical Information Retrieval without Relevance Labels

### v1

@@ 5.3. Performance Analysis
Effect with Different Retrievers   We consider fine-tuning the other two retrieval models: PEG which achieves optimal performance on CMIRB, and a multilingual retrieval mE5.
@@
How to Use Hypothetical Documents   In this section, we test several methods for incorporating hypothetical documents. SL-HyDE: This method encodes the original query and the hypothetical documents separately, then applies mean pooling to obtain the final query vector. SL-HyDE w/K-D: This approach generates five hypothetical documents. SL-HyDE w/D: Only the hypothetical document is used as the query for retrieval. SL-HyDE w/con: The original query and the hypothetical document are concatenated into a single string to form a new query.
Table 5 shows that the combination of the original query and hypothetical documents is optimal. Sole reliance on hypothetical documents significantly reduces performance, especially in medical consultation tasks, where original queries contain critical information. The string concatenation method introduces some performance degradation, indicating that the generated documents may contain noise at the string level, whereas average pooling effectively mitigates it. Generating multiple hypothetical documents increases coverage and improves performance across tasks. However, it often leads to a K-fold increase in inference time.
Ablation Study    To further analyze the gains brought by the internal architecture of the SL-HyDE model, we conduct two sets of ablation experiments. (1) A variant, SL-HyDE w/o BGE-FT, which uses the fine-tuned LLM as the generator and the raw BGE as the retriever; (2) A variant, SL-HyDE w/o Qwen-FT, which utilizes raw LLM as the generator and the fine-tuned BGE as the retriever.
Table 6 demonstrates that fine-tuning both components substantially enhances performance, validating the efficacy of the self-learning mechanism. Notably, fine-tuning the retriever yields greater gains, suggesting that BGE benefits significantly from domain-specific adaptation. However, our approach does not rely on labeled data and still achieves significant retrieval performance improvements.
Table 5. Performance of different fusing strategies.
@@
Table 6. Performance of different variants.
@@
6. CONCLUSIONS
@@
In this paper, we introduce an automated framework for zero-shot medical information retrieval, named SL-HyDE, which operates without the need for relevance labels. Utilizing an unlabeled medical corpus, we employ a self-learning, end-to-end training framework where the retriever guides the generator’s training, and the generator, in turn, enhances the retriever. This process integrates medical knowledge to create hypothetical documents that are more effective in retrieving target documents. Furthermore, we present a comprehensive Chinese medical information retrieval benchmark, evaluating mainstream retrieval models against this new standard. Experimental findings demonstrate that SL-HyDE consistently improves retrieval accuracy over HyDE across ten datasets. Additionally, SL-HyDE shows strong adaptability and scalability, effectively enhancing retrieval performance across various combinations of generators and retrievers.
Appendix A CMIRB DATASETS
A.1. Data Process
We curated a substantial dataset from various medical resources, as presented in Table 7, which details the source distribution and data volume. Our data preprocessing pipeline, depicted in Algorithm 1, employs prompt templates outlined in Tables 11 and 12.
Initially, we use ChatGPT [^11] https://openai.com/chatgpt to perform medical relevance detection on the texts, eliminating non-medical content (lines 3-8). Subsequently, ChatGPT assesses query-document relevance, filtering out low-relevance examples (lines 27-33). Our relevance assessment considers semantic alignment and the practical significance of data samples for their respective tasks, as highlighted in prompt 12.
For the MedExam and DuBaike datasets, the direct query-document signal isn’t initially provided. Both queries and documents in the MedExam dataset originate from Work (Jin et al., 2021), where 100 randomly selected questions have corpus documents containing evidence sufficient to answer them, verified manually by the authors. In the DuBaike dataset, queries from Baidu Search and Baidu Zhidao often match the content distribution of Baidu Baike. These factors allow us to design a query-matching algorithm to locate the valuable document.
Through this iterative loop of self-ranking, evidence searching, answering, and verification, combined with ChatGPT’s advanced knowledge capabilities, we ensure high-quality, highly relevant query-document pairs. Post-processing statistical details of these datasets are displayed in Table 1.
Algorithm 1 Data Preprocessing Pipeline
3:// Step 1: Filter out medically irrelevant queries and documents
4:for each query $q\in Q$, $d\in D$ do
5:     $med_{score}\leftarrow\mathrm{LLM}.\text{med\_score}(q/d)$
6:     if $med_{score}<threshold$ then
7:         Remove $q/d$
8:     end if
9:end for
10:// Step 2: Matching positive query-document pair
11:if query-document matching is required then
12:     for each query $q\in Q$ do
13:         // Retrieve top-k relevant documents
14:         $d_{1},d_{2},\ldots,d_{M}\leftarrow\text{BM25}(q,D)$
15:         $d_{1},d_{2},\ldots,d_{k}\leftarrow\mathrm{LLM}.\text{reranking}(q;d_{1},d_{2},\ldots,d_{M})$
16:         // Extract evidence snippets
17:         $e_{1},e_{2},\ldots,e_{k}\leftarrow\mathrm{LLM}.\text{extract\_evidence}(q,d_{1}|d_{2}|\ldots|d_{k})$
18:         // Generate answers
19:         $a_{1},a_{2},\ldots,a_{k}\leftarrow\mathrm{LLM}.\text{answer}(q,e_{1}|e_{2}|\ldots|e_{k})$
20:         for each document $d_{i}$ do
21:              if $\mathrm{LLM}.\text{validate}(a_{i},d_{i})$ is correct then
22:                  Store $(q,d_{i})$ as a matched pair
23:              end if
24:         end for
25:     end for
26:end if
27:// Step 3: Filter out pseudo-relevant match pairs
28:for each matched pair $(q,d)$ do
29:     $rel_{score}\leftarrow\mathrm{LLM}.\text{filter\_score}(q,d)$
30:     if $rel_{score}<threshold$ then
31:         Remove $(q,d)$
32:     end if
33:end for
Table 7. Dataset Collection Sources and Quantity Statistics.
A.2. Data Example
The datasets we constructed encompass various real-world medical scenarios, with examples from 10 different datasets illustrated in Figure 3. Queries can take the form of a medical paper title, a patient’s symptom description, or an exam question. Corresponding documents include abstracts of medical papers, doctor-patient diagnostic conversations, and reference materials for exam questions.
Appendix B MODELS
B.1. Baselines
@@
To comprehensively evaluate the performance of existing retrievers on CMIRB, we selected 10 representative models, all of which have achieved strong results on the MTEB leaderboard [^12] https://huggingface.co/spaces/mteb/leaderboard.
@@
BM25.   BM25 (Robertson et al., 2009) is a commonly used baseline retriever which uses bag-of-words and TF-IDF to perform lexical retrieval. In this paper, BM25 is implemented with Pyserini (Lin et al., 2021) using the default hyperparameters to index snippets from all corpora.
@@
Text2Vec [^13] https://huggingface.co/GanymedeNil/text2vec-large-chinese.   It is a cosine sentence model based on a linguistically-motivated pre-trained language model (LERT).
@@
PEG [^14] https://huggingface.co/TownsWu/PEG.    Wu et al., (Wu et al., 2023) proposes the PEG, which is trained on more than 100 million data, encompassing a wide range of domains and covering various tasks.
@@
BGE [^15] https://huggingface.co/BAAI/bge-large-zh-v1.5.    It takes a compound recipe to train general-purpose text embedding, including, embedding-oriented pre-training, contrastive learning with sophisticated negative sampling, and instruction-based fine-tuning.
@@
GTE [^16] https://huggingface.co/thenlper/gte-large-zh.    It presents a multi-stage contrastive learning approach to develop text embedding model that can be applied to various tasks.
@@
Piccolo [^17] https://huggingface.co/sensenova/piccolo-large-zh.    Piccolo is a general-purpose Chinese embedding model trained using a two-stage process with weakly supervised and manually labeled text pairs.
@@
Contriever  [^18] https://huggingface.co/facebook/mcontriever-msmarco.    It is a multilingual dense retriever with contrastive learning, which fine-tunes the pre-trained mContriever model on MS MARCO dataset.
@@
M3E  [^19] https://huggingface.co/moka-ai/m3e-large.    M3E (Moka Massive Mixed Embedding) is a bilingual text embedding model trained on over 22 million Chinese sentence pairs, supporting tasks like cross-lingual text similarity and retrieval.
Table 8. Evaluation Prompts for LLMs.
@@
Please generate a medical content paragraph to answer this question.
Question: [QUESTION]
@@
Please generate a medical content paragraph based on this title.
Title: [TITLE]
@@
Please generate a similar medical paragraph for the following text.
Text: [TEXT]
@@
mE5  [^20] https://huggingface.co/intfloat/multilingual-e5-large.   Multilingual E5 text embedding models that are trained with a multi-stage pipeline, involving contrastive pre-training on 1 billion multilingual text pairs, and fine-tuning on labeled datasets.
Table 9. Performance of various Retrieval models on CMIRB benchmark. All scores denote Recall@100. The best score on a given dataset is marked in bold.
@@
Table 10. Performance of different combinations of generators and retrievers on CMIRB benchmark.

### v2

@@ 5.3 Performance Analysis
Table 3: Performance of different generators.
Effect of Different Retrievers. We consider fine-tuning the other two retrieval models: PEG which achieves optimal performance on CMIRB, and a multilingual retriever mE5.
@@
Table 4: Performance of different retrievers.
@@
Table 5: Performance of different fusing strategies.
Effect of Different Fusing Strategies. In this section, we test several methods for incorporating hypothetical documents. SL-HyDE: This method encodes the original query and the hypothetical documents separately, then applies mean pooling to obtain the final query vector. SL-HyDE w/ D: Only the hypothetical document is used as the query for retrieval. SL-HyDE w/ con: The original query and the hypothetical document are concatenated into a single string to form a new query. SL-HyDE w/ K-D: This approach generates five documents.
Table 5 shows that the combination of the original query and hypothetical documents is optimal. Sole reliance on hypothetical documents significantly reduces performance, especially in medical consultation tasks, where original queries contain critical information. The string concatenation method introduces some performance degradation, indicating that the generated documents may contain noise at the string level, whereas average pooling effectively mitigates it. Generating multiple hypothetical documents increases coverage and improves performance across tasks. However, it often leads to a K-fold increase in inference time. Therefore, we need to balance efficiency and accuracy to select the number of hypothetical documents.
5.4 Ablation Study
To further analyze the gains brought by the internal architecture of SL-HyDE, we conduct two sets of ablation experiments: (1) SL-HyDE w/o BGE-FT, which uses the fine-tuned LLM as the generator and the raw BGE as the retriever; (2) SL-HyDE w/o Qwen-FT, which utilizes raw LLM as the generator and the fine-tuned BGE as the retriever.
Table 6 demonstrates that fine-tuning both components substantially enhances performance, validating the efficacy of the self-learning mechanism. Notably, fine-tuning the retriever yields greater gains, suggesting that BGE benefits significantly from domain-specific adaptation. However, our approach fine-tunes both the retriever and the generator, boosting their performance between the two to enhance retrieval tasks.
@@ 5.4 Ablation Study
Table 6: Performance of different variants.
5.5 Case Study
To intuitively show how the SL-HyDE makes a difference in the hypothetical documents and retrieval performance, we present examples in Table 7 to compare the hypothetical document generated by HyDE and SL-HyDE. The query is How to treat a hernia?. While HyDE generates a general document discussing conservative and surgical treatments, it lacks specificity for different patient groups. In contrast, SL-HyDE produces a document mentioning hernia belts for infants and elderly patients, closely matching the target document’s details. This improved relevance led to a higher retrieval ranking ($2nd$ vs. $10th$), demonstrating how more precise hypothetical documents enhance retrieval performance.
Query: How to treat a hernia?
Target Doc: Inguinal Hernia Treatment Plan. For conventional treatment,
a 1-year-old infant can use a hernia belt for compression. As the muscles
gradually strengthen, there may be a possibility of spontaneous recovery.
For elderly and frail individuals a hernia belt can be worn, but for
other patients, surgery is generally recommended…
HyDE: Hernia is a common disease caused by a weak area in the
abdominal wall, Treatment usually includes conservative and surgical
methods. For most patients, especially young and healthy individuals,
surgery is the preferred option… (Rank: 10)
SL-HyDE: Hernia is a common condition that typically occurs… For
infants,… the use of a hernia belt to apply localized pressure can help
alleviate symptoms and promote the development of the abdominal
muscles,… For elderly or frail patients, or those with severe underlying
conditions,… wearing a hernia belt can help manage symptoms and
reduce the risk of the hernia progressing further… (Rank: 2)
Table 7: The case study comparing with baseline.
6 Conclusions
@@
In this paper, we introduce an automated framework for zero-shot medical information retrieval, named SL-HyDE, which operates without the need for relevance labels. Utilizing an unlabeled medical corpus, we employ a self-learning, end-to-end training framework where the retriever guides the generator’s training, and the generator, in turn, enhances the retriever. This process integrates medical knowledge to create hypothetical documents that are more effective in retrieving target documents. Furthermore, we present a comprehensive Chinese medical information retrieval benchmark, evaluating mainstream retrieval models against this new standard. Experimental findings demonstrate that SL-HyDE consistently improves retrieval accuracy over HyDE across ten datasets. Additionally, SL-HyDE shows strong adaptability and scalability, effectively enhancing retrieval performance across various combinations of generators and retrievers. In future work, we will extend SL-HyDE to other data-scarce domains to further evaluate its generalizability across different settings. In addition, we will explore reinforcement learning to train more capable retrievers and enhance reasoning in complex medical retrieval tasks.
7 Limitations
While our work effectively addresses the adaptation challenges of HyDE in low-resource scenarios, several limitations remain. First, our study primarily focuses on the medical domain and provides a preliminary exploration in the legal domain (see Appendix A.4), but we have not extended our investigation to other vertical domains such as economics or education. Second, although we experiment with three open-source LLMs, Qwen2, LlaMA2, and ChatGLM3, as generators, we do not include more recent or diverse model families such as Qwen3 or Gemini, which may exhibit different generation behaviors. Third, our data construction pipeline relies on LLMs for query-document matching and pseudo-relevant pair filtering. The effectiveness of these components depends on the model’s instruction-following ability and its sensitivity to domain-specific nuances, which may introduce hallucinations or spurious correlations.
Appendix A Models
A.1 Baselines
@@
To comprehensively evaluate the performance of existing retrievers on CMIRB, we selected 10 representative models, all of which have achieved strong results on the MTEB leaderboard [^5] https://huggingface.co/spaces/mteb/leaderboard. For details regarding the retrievers and large reasoning models evaluated throughout the paper, please refer to Table 8.
@@
BM25 Robertson et al. (2009).   BM25 is a commonly used baseline retriever which uses bag-of-words and TF-IDF to perform lexical retrieval. In this paper, BM25 is implemented with Pyserini Lin et al. (2021) using the default hyperparameters to index snippets from all corpora.
@@
Text2Vec Xu (2023).   It is a cosine sentence model based on a linguistically-motivated pre-trained language model (LERT).
@@
PEG Wu et al. (2023).    Wu et al., Wu et al. (2023) proposes the PEG, which is trained on more than 100 million data, encompassing a wide range of domains and covering various tasks.
@@
BGE Xiao et al. (2024).    It takes a compound recipe to train general-purpose text embedding, including, embedding-oriented pre-training, contrastive learning with sophisticated negative sampling, and instruction-based fine-tuning.
@@
GTE Li et al. (2023b).    It presents a multi-stage contrastive learning approach to develop text embedding model that can be applied to various tasks.
@@
Piccolo SenseTime (2023).    Piccolo is a general-purpose Chinese embedding model trained using a two-stage process with weakly supervised and manually labeled text pairs.
@@
Contriever  Izacard et al. (2021).    It is a multilingual dense retriever with contrastive learning, which fine-tunes the pre-trained mContriever model on MS MARCO dataset.
@@
M3E Wang et al. (2023b).    M3E (Moka Massive Mixed Embedding) is a bilingual text embedding model trained on over 22 million Chinese sentence pairs, supporting tasks like cross-lingual text similarity and retrieval.
@@
Please generate a medical content paragraph to answer
this question.
Question: QUESTION
@@
Please generate a medical content paragraph based on
this title.
Title: TITLE
@@
Please generate a similar medical paragraph for the
following text.
Text: TEXT
@@
Table 9: Evaluation prompts for generators.
mE5 Wang et al. (2024a).   Multilingual E5 text embedding models that are trained with a multi-stage pipeline, involving contrastive pre-training on 1 billion multilingual text pairs, and fine-tuning on labeled datasets.
OpenAI-Ada-002 OpenAI .   It is a highly efficient text embedding model that converts natural language into dense vectors for a wide range of applications, including semantic search and similarity tasks.
Qwen2 Yang et al. (2024).   Qwen2 is a comprehensive suite of foundational and instruction-tuned language models, encompassing a parameter range from 0.5 to 72 billion, featuring dense models and a Mixture-of-Experts model.
ChatGLM3 Team et al. (2024).   ChatGLM3-6B is a next-generation conversational pre-trained model with strong performance across tasks like semantics, reasoning, and code execution, and supports complex scenarios such as tool use and function calls.
Llama2 Touvron et al. (2023).    Llama2 is an auto-regressive language model that uses an optimized transformer architecture. The tuned versions utilize supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align with human preferences for helpfulness and safety.
A.2 Evaluation Settings
We use the C-MTEB [^6] C-MTEB framework to evaluate the performance of various retrieval models on CMIRB. To ensure stability, we set the temperature of LLM to 0.7 and repeat five times with different random seeds. For each dataset, the prompts used to generate pseudo-documents are shown in Figure 9. The IIYIPost and CSLCite datasets utilize the T2P template to prompt LLMs to generate documents based on the given title. For the CSLRel dataset, we employ the P2P template to instruct the model to produce similar text. As for the other datasets, the Q2P template is employed by the LLM to generate answers to medical questions.
A.3 SL-HyDE vs. HyDE
Our approach, SL-HyDE, builds upon HyDE Gao et al. (2022) with several enhancements while retaining some similarities. First, both SL-HyDE and HyDE follow the same inference process. Each uses a large model to generate a hypothetical document based on the query, which the retriever then employs to locate the most relevant document. Second, neither SL-HyDE nor HyDE requires labeled data, which allows for rapid deployment. HyDE is especially advantageous in real-world scenarios where efficient retrieval can be executed simply by selecting a generator and a retriever. However, for tasks needing domain-specific knowledge, such as medical information retrieval, deploying HyDE directly may not yield optimal results. One potential strategy is to fine-tune the generator and retriever separately using labeled medical data before deploying the HyDE framework. The primary challenge here in acquiring labeled data, and fine-tuning the models separately often leads to suboptimal performance.
SL-HyDE improves upon this by integrating a self-learning mechanism, transforming HyDE into a trainable end-to-end framework. This mechanism enables both the generator and the retriever to better adapt to the medical domain. Supervision signals for the generator’s training are derived from the retriever, and vice versa, facilitating mutual enhancement through this self-learning process. This holistic approach results in improved performance in retrieval tasks. Overall, SL-HyDE offers an efficient and convenient solution for enhancing HyDE’s performance in the medical domain, particularly when dealing with unlabeled corpora.
A.4 More Experiment Results
SL-HyDE can be easily applied to other domains that lack labeled data. By fine-tuning both the generator and retriever using only a small amount of unstructured domain text, it builds an effective retrieval system. Specifically, we apply SL-HyDE to the English legal domain. We sample 10k law texts from pile-of-law  [^7] https://huggingface.co/datasets/pile-of-law/pile-of-law and use Llama-2-7b-chat-hf as the generator and BGE-Large-en-V1.5 as the retriever. We evaluate three information retrieval datasets in the law domain from MTEB. The results in Table 12 shows that SL-HyDE (77.25%) significantly outperforms HyDE (75.52%) in the legal domain.
@@ A.4 More Experiment Results
Table 10: Performance of various Retrieval models on CMIRB benchmark. All scores denote Recall@100. The best score on a given dataset is marked in bold.